# Logarithmic Enriques varieties

### v1

@@ 6.2.2. Log-Enriques varieties simultaneously of IHS and CY type
The induced automorphisms $\iota,\sigma\in\mathrm{Aut}(F(C))$ act nonsymplectically on the Fano variety of lines $F(C)$ (see [18, §7] and [12, Example 6.4] for details). The quotient $Y:=F(C)/\iota$ is a singular Calabi–Yau variety, as discussed in [22], and the order three automorphism $\sigma$ descends to an automorphism $\overline{\sigma}$ of the quotient $Y$; moreover, $\overline{\sigma}$ does not preserve the volume form on $Y$. Hence, the quotient $X:=Y/\overline{\sigma}=F(C)/\iota\circ\sigma$ is a log-Enriques variety of index 3, both of CY type and of IHS type.
@@ 6.2.3. Index $6$ log-Enriques varieties
Let $A=E\times E_{\omega}$ where $E$ is an elliptic curve, and $E_{\omega}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\omega\mathbb{Z}}$. Take $u\in E$, $v\in E_{\omega}$ and consider the order $6$ automorphism $f_{A}(x,y)=(-x+u,\omega y+v)$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$. By Proposition 5.4, we get that the $2n$-dimensional quotient ${\rm K}_{n}(A)/{\rm K}_{n}(f_{A})$ has trivial canonical bundle if $n\equiv 0\mod 6$, but it is a log-Enriques variety of IHS type otherwise:
@@ 6.2.4. Index $12$ log-Enriques varieties
Let $A=E_{\mathrm{i}}\times E_{\omega}$ where $\omega$ is a primitive $3$rd root of unity, $E_{\mathrm{i}}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\mathrm{i}\mathbb{Z}}$, $\omega$ and $E_{\omega}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\omega\mathbb{Z}}$. Take $u\in E_{\mathrm{i}}$, $v\in E_{\omega}$ and consider the order $12$ automorphism $f_{A}(x,y)=(\mathrm{i}x+u,\omega y+v)$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$. By Proposition 5.4, we get that the $2n$-dimensional quotient ${\rm K}_{n}(A)/{\rm K}_{n}(f_{A})$ has trivial canonical bundle if $n\equiv 0\mod 12$, but it is a log-Enriques variety of IHS type otherwise:
@@ 6.3.1. Index $2$ log-Enriques varieties from moduli spaces of semistable sheaves on $K3$ surfaces
Here, we observe that Oguiso–Schröer’s construction of Enriques manifolds illustrated in §6.1.4 generalizes to singular moduli spaces of semistable sheaves on $K3$ surfaces, which are known to be ISV [50]. Indeed, let $S$ be a very general $K3$ surface carrying an Enriques involution $\iota_{S}$. By genericity we can assume that $\rho(S)=10$ and that the Picard lattice $\mathrm{Pic}(S)$ is $\iota_{S}$-invariant. Given a non-primitive Mukai vector $w=(r,l,\chi-r)\in H^{*}(S,\mathbb{Z})$ such that $w^{2}\equiv 0\mod 4$ and a $w$-general polarization $H\in\mathrm{Pic}(S)$, the induced involution $\iota_{S}^{*}$ acts on the moduli space $M_{w,H}(S)$ of $H$-semistable sheaves on $S$ with Mukai vector $w$, and the quotient $M_{w,H}(S)/\iota_{S}^{*}$ is a log-Enriques variety of dimension $w^{2}+2$, of ISV type and index two.
Moreover, if $\chi$ is odd, $\iota_{S}^{*}$ acts freely on the regular locus of $M_{w,H}(S)$, which is the locus of $H$-stable sheaves on $S$. In order to show this, it is enough to observe that the proof of  [47, Theorem 5.3] works: if a stable sheaf $\mathcal{F}$ is fixed, by descent there exists a coherent sheaf $\mathcal{F}^{\prime}$ on $S/\iota_{S}$ such that $\mathcal{F}=p^{*}\mathcal{F}^{\prime}$, and thus $\chi=\chi(\mathcal{F})=2\chi(\mathcal{F}^{\prime})$, in contradiction with $\chi$ being odd.
@@ 6.3.2. Index $2$ log-Enriques varieties from relative Prym varieties
Let $S$ be a very general $K3$ surface carrying an Enriques involution $i_{S}$, as in §6.1.4 and §6.3.1. Let $\pi\colon S\to T:=S/i_{S}$ be the quotient Enriques surface, let $C$ be a smooth curve of genus $g$ on $T$ with primitive class in $\mathrm{NS}(T)$ and denote $D:=\pi^{-1}(C)$. The moduli space $M_{v,D}(S)$ is singular when $v=(0,D,2-2g)\in H^{*}(S,\mathbb{Z})$ (see  [2]) and it is an irreducible symplectic variety (see [16] and references therein). Moreover, since $D^{2}=4g-4$, it is of dimension $2(2g-1)$, hence it follows from §6.3.1 that $i^{*}_{S}$ acts freely on $(M_{v,D}(S))_{\rm reg}$. On the other hand, the moduli space is also endowed with the duality $j$, which is a non-symplectic involution commuting with $i^{*}_{S}$. The connected component $\mathcal{P}_{v,D}$ of the zero-section inside the fixed locus of the symplectic involution $i^{*}_{S}\circ j$ is the so-called relative Prym variety, and it is an irreducible symplectic variety by [2]. The non-symplectic involution $i^{*}_{S}$ restricts to $\mathcal{P}_{v,D}$ and acts freely on $\mathcal{P}_{v,D}\cap(M_{v,D}(S))_{\rm reg}$. For all genera $g\geq 2$ then $\mathcal{P}_{v,D}/i^{*}_{S}$ is a log-Enriques variety of ISV type and of index two.
This construction also works when $S$ is a very general $K3$ surface endowed with a non-symplectic involution $i$ such that $\mathrm{NS}(S)\simeq U(2)\oplus E_{8}(-1)^{\oplus 2}$ or $\mathrm{NS}(S)\simeq U\oplus E_{8}(-1)\oplus D_{4}(-1)^{\oplus 2}$. In such a case, the quotient surface $T$ is rational. By [16], one can also construct relative Prym varieties $\mathcal{P}_{v,D}$ starting from a curve $C$ on $T$ of genus $g$ and considering the moduli space $M_{v,D}(S)$ of $D$-semistable sheaves on $S$ with Mukai vector $(0,D,1-g(D))$, where $D:=\pi^{-1}(C)$ is of genus $g(D)$. Moreover, in this case $S$ also carries an Enriques involution $i_{S}$ by [48], which commutes with $i$. As a consequence, whenever $D$ is also invariant for $i_{S}$, the involution $i_{S}^{*}$ acts on $\mathcal{P}_{v,D}$ and the quotient $\mathcal{P}_{v,D}/i^{*}_{S}$ is a log-Enriques variety of index two, which is of PSV type (resp. of ISV type) when $C$ and $D$ satisfy the assumptions of [16, Theorem 1.2] (resp. of [16, Theorem 1.3 and 1.4]).

### v2

@@ 6.2.2. Log-Enriques varieties of IHS type from weak CY varieties
The induced automorphisms $\iota,\sigma\in\mathrm{Aut}(F(C))$ act nonsymplectically on the Fano variety of lines $F(C)$ (see [20, §7] and [14, Example 6.4]). The quotient $Y:=F(C)/\langle\iota\rangle$ is a singular weak Calabi–Yau variety [24] and the order three automorphism $\sigma$ descends to an automorphism $\overline{\sigma}$ of the quotient $Y$; moreover, $\overline{\sigma}$ does not preserve the volume form on $Y$. Hence, the quotient:
$X:=Y/\langle\overline{\sigma}\rangle=F(C)/\langle\iota\circ\sigma\rangle,$
is a log-Enriques variety of IHS type and index three. On the other hand one can consider first the quotient $F(C)/\langle\sigma\rangle$. The fixed locus of $\sigma$ on $F(C)$ is a smooth surface (see [14, Example 6.4]), its image will determine the singular locus on $F(C)/\langle\sigma\rangle$. By using the formula of §6.2.1 we compute the age at a singular point to be equal to $2/3$, this means that the singularities are here strictly klt. Now the quotient $F(C)/\langle\iota\circ\sigma\rangle$ has klt singularities by [44, Proposition 5.20] and since by $loc.\,cit.$ discrepancies decrease by doing finite quotients we get that $X$ is a uniruled log-Enriques variety.
@@ 6.2.3. Index $6$ log-Enriques varieties
Let $A=E\times E_{\omega}$ where $E$ is an elliptic curve, and $E_{\omega}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\omega\mathbb{Z}}$ where $\omega$ is a primitive third root of unity. Take $u\in E$, $v\in E_{\omega}$ and consider the order $6$ automorphism $f_{A}(x,y)=(-x+u,\omega y+v)$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$. By Proposition 5.4, we get that the $2n$-dimensional quotient ${\rm K}_{n}(A)/\langle{\rm K}_{n}(f_{A})\rangle$ has trivial canonical bundle if $n\equiv 0\mod 6$, but it is a log-Enriques variety of IHS type otherwise:
@@ 6.2.4. Index $12$ log-Enriques varieties
Let $A=E_{\mathrm{i}}\times E_{\omega}$ where $\omega$ is a primitive $3$rd root of unity, $E_{\mathrm{i}}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\mathrm{i}\mathbb{Z}}$ and $E_{\omega}=\frac{\mathbb{C}}{\mathbb{Z}\oplus\omega\mathbb{Z}}$. Take $u\in E_{\mathrm{i}}$, $v\in E_{\omega}$ and consider the order $12$ automorphism $f_{A}(x,y)=(\mathrm{i}x+u,\omega y+v)$. Assuming that $u$ and $v$ are $(n+1)$-torsion points, the natural automorphism ${f_{A}}^{[n+1]}$ of ${A}^{[n+1]}$ respects the fiber ${\rm K}_{n}(A)$. By Proposition 5.4, we get that the $2n$-dimensional quotient ${\rm K}_{n}(A)/\langle{\rm K}_{n}(f_{A})\rangle$ has trivial canonical bundle if $n\equiv 0\mod 12$, but it is a log-Enriques variety of IHS type otherwise:
@@ 6.3.1. Index $2$ log-Enriques varieties from moduli spaces of semistable sheaves on $K3$ surfaces
Here, we observe that Oguiso–Schröer’s construction of Enriques manifolds illustrated in §6.1.4 generalizes to singular moduli spaces of semistable sheaves on $K3$ surfaces, which are known to be ISV [55]. Indeed, let $S$ be a very general $K3$ surface carrying an Enriques involution $\iota_{S}$. By genericity we can assume that $\rho(S)=10$ and that the Picard lattice $\mathrm{Pic}(S)$ is $\iota_{S}$-invariant. Given a non-primitive Mukai vector $w=(r,l,\chi-r)\in H^{*}(S,\mathbb{Z})$ such that $w^{2}\equiv 0\mod 4$ and a $w$-general polarization $H\in\mathrm{Pic}(S)$, the induced involution $\iota_{S}^{*}$ acts on the moduli space $M_{w,H}(S)$ of $H$-semistable sheaves on $S$ with Mukai vector $w$, and the quotient $M_{w,H}(S)/\langle\iota_{S}^{*}\rangle$ is a log-Enriques variety of dimension $w^{2}+2$, of ISV type and index two.
Moreover, if $\chi$ is odd, $\iota_{S}^{*}$ acts freely on the regular locus of $M_{w,H}(S)$, which is the locus of $H$-stable sheaves on $S$. In order to show this, it is enough to observe that the proof of  [51, Theorem 5.3] works: if a stable sheaf $\mathcal{F}$ is fixed, by descent there exists a coherent sheaf $\mathcal{F}^{\prime}$ on $S/\langle\iota_{S}\rangle$ such that $\mathcal{F}=p^{*}\mathcal{F}^{\prime}$, and thus $\chi=\chi(\mathcal{F})=2\chi(\mathcal{F}^{\prime})$, in contradiction with $\chi$ being odd.
@@ 6.3.2. Index $2$ log-Enriques varieties from relative Prym varieties
Let $S$ be a very general $K3$ surface carrying an Enriques involution $i_{S}$, as in §6.1.4 and §6.3.1. Let $\pi\colon S\to T:=S/\langle i_{S}\rangle$ be the quotient Enriques surface, let $C$ be a smooth curve of genus $g$ on $T$ with primitive class in $\mathrm{NS}(T)$ and denote $D:=\pi^{-1}(C)$. The moduli space $M_{v,D}(S)$ is singular when $v=(0,D,2-2g)\in H^{*}(S,\mathbb{Z})$ (see  [2]) and it is an irreducible symplectic variety (see [18] and references therein). Moreover, since $D^{2}=4g-4$, it is of dimension $2(2g-1)$, hence it follows from §6.3.1 that $i^{*}_{S}$ acts freely on $(M_{v,D}(S))_{\rm reg}$. On the other hand, the moduli space is also endowed with the duality $j$, which is a nonsymplectic involution commuting with $i^{*}_{S}$. The connected component $\mathcal{P}_{v,D}$ of the zero-section inside the fixed locus of the symplectic involution $i^{*}_{S}\circ j$ is the so-called relative Prym variety, and it is an irreducible symplectic variety by [2]. The nonsymplectic involution $i^{*}_{S}$ restricts to $\mathcal{P}_{v,D}$ and acts freely on $\mathcal{P}_{v,D}\cap(M_{v,D}(S))_{\rm reg}$. For all genera $g\geq 2$ then $\mathcal{P}_{v,D}/i^{*}_{S}$ is a log-Enriques variety of ISV type and of index two.
This construction also works when $S$ is a very general $K3$ surface endowed with a nonsymplectic involution $i$ such that $\mathrm{NS}(S)\simeq U(2)\oplus E_{8}(-1)^{\oplus 2}$ or $\mathrm{NS}(S)\simeq U\oplus E_{8}(-1)\oplus D_{4}(-1)^{\oplus 2}$. In such a case, the quotient surface $T$ is rational. By [18], one can also construct relative Prym varieties $\mathcal{P}_{v,D}$ starting from a curve $C$ on $T$ of genus $g$ and considering the moduli space $M_{v,D}(S)$ of $D$-semistable sheaves on $S$ with Mukai vector $(0,D,1-g(D))$, where $D:=\pi^{-1}(C)$ is of genus $g(D)$. Moreover, in this case $S$ also carries an Enriques involution $i_{S}$ by [52], which commutes with $i$. As a consequence, whenever $D$ is also invariant for $i_{S}$, the involution $i_{S}^{*}$ acts on $\mathcal{P}_{v,D}$ and the quotient $\mathcal{P}_{v,D}/\langle i^{*}_{S}\rangle$ is a log-Enriques variety of index two, which is of PSV type (resp. of ISV type) when $C$ and $D$ satisfy the assumptions of [18, Theorem 1.2] (resp. of [18, Theorem 1.3 and 1.4]).